First of all, I would like to take this opportunity to thank you for taking the time to answer my question. Here it goes:
In class, we are studying how to solve systems of linear equations using substitution. The answer key I was given said that this question was supposed to have the solution set of {(-10,2)} The answer I figured out was {(70,-82})
The question was 3x+2y= -34
y-x= 12
I would really appreciate a reply, for I would like to know how to do it before we get to the next step of solving linear equations. Thank you again for your reply on this matter.

Answer by rapaljer(4671)

Use the second equation to solve for y by adding x to each side:
y-x = 12
y=x+ 12

Now substitute into the first equation:
3x + 2y = -34
3x + 2(_______) = -34
3x + 2(x+ 12)= -34
3x + 2x + 24 = -34
5x +24 - 24 = -34 - 24
5x = -58
x= -58/5

I think maybe you copied the problem wrong. Better take another look. I think the second equation was supposed to be 3x-2y = -34. If this is the case, then the answer in the solution was correct, and you made an error.

R^2 at SCC